The coordinate of a particle in meters is given by x(t)=2.0t-2.0t², where the time t is in seconds. The particle is momentarily at rest at time t equal to?
1
Expert's answer
2020-10-16T10:57:55-0400

If the particle is at rest, then its velocity is zero, so we find the velocity equation, which is equal to the derivative of the coordinate equation "v(t)=(x(t))^{\prime}=2-4\times t" , for "v=0" it will be

"0=2-4\times t"

"t=0.5" s.